28, Mo EO, o=0 o=0 X 45° Hi o k

medium1 : ε0, μ0

medium2 : 2ε0, μ0

Condition : 1GHz TM electromagnetic wave, Hi0 = 1A/m

What is poynting vector of field that exists in medium2?

I think that we can't calculate the ponyting vector of field exists in medium 2 in the total reflection case.
Answer #1

Answer :

Given data :

Medium 1 : ε0, μ0

medium 2 : 2ε0, μ0

angle \theta = 45\degree

So Know that,

​​Poynting vector \vec{S} = \vec{E}\times \vec{H}    and

S = EH \sin\theta

where E = \mu_o c H

So, \bf S = \mu_o c H^2\sin \theta    = S =\frac{\mu_0c(1)^2}{\sqrt2} = S =\frac{\mu_0c}{\sqrt2}    .

It is a case of total reflection,

because in this case,

refractive index of medium 2 = √2 n

refractive index of medium 1 = n

by applying snell's law,

we can find the value of angle of refraction throgh medium 1 = 90° which is a total reflection case.